OpenStudy (anonymous):

Let A be the area of a circle with radius r. If dr/dt = 3, find dA/dt when r = 1

OpenStudy (vishal_kothari):



sam (.sam.):

\[\huge \frac{dA}{dt}= \frac{dA}{dr}\frac{dr}{dt}\]

sam (.sam.):

chain rule

sam (.sam.):

\[\huge A=\pi r ^{2}\] \[\huge \frac{dA}{dr}=2\pi r \] When r=1 \[\huge \frac{dA}{dr}=2\pi \]

sam (.sam.):

\[\huge \frac{dA}{dt}=2 \pi (3)=6\pi\]